Consider the following historical rates of return: Year Stock X returns

Consider the following historical rates of return:

                         Year   Stock X returns     

                        2014 -0.12   

                        2015                0.04            

                        2016               -0.01             

                        2017 0.05

                        2018               0.10

2019 0.08

a.  Find standard deviation of these returns.

b. Find coefficient of variation (CV).

Please show work how to find each!
Answer #1

Part a

Standard deviation is a measure of Total risk

Standard Deviation = \sqrt{\frac{\sum (Given Return -Average Return)^{2}}{number of years}}

Year Return (Return - Average Return)2
Y X (X - X')2
2014 -12 164.6944
2015 4 10.0278
2016 -1 3.3611
2017 5 17.3611
2018 1 0.0278
2019 8 51.3611
Total ( a ) 5 246.8333

Average Return = Total Returns / number of years = 5 / 6 = 0.8333 %

Standard deviation =\sqrt{\frac{246.8333}{6}}

Standard deviation = 6.414

Part B

Coefficient of Variation (CV) is defined as "Risk per unit of return"

Coefficient of Variation (CV) = Standard Deviation \div Average Return

= 6.414 \div 0.8333

CV = 7.697

Please upvote the answer if it helped you.
In case of any doubt just comment below, I would love to help.